The American Heart Association is about to conduct an anti-smoking campaign and wants to know the fraction of Americans over 4444 who smoke.

Step 1 of 2:

Suppose a sample of 632632 Americans over 4444 is drawn. Of these people, 170170 smoke. Using the data, estimate the proportion of Americans over 4444 who smoke. Enter your answer as a fraction or a decimal number rounded to three decimal places.

The American Heart Association is about to conduct an anti-smoking campaign and wants to know the fraction of Americans over 4444 who smoke.

Step 2 of 2:

Suppose a sample of 632632 Americans over 4444 is drawn. Of these people, 170170 smoke. Using the data, construct the 90%90% confidence interval for the population proportion of Americans over 4444 who smoke. Round your answers to three decimal places.
Answer #1

1)
sample proportion, = 170/632 = 0.269


b)


sample size, n = 632
Standard error, SE = sqrt(pcap * (1 - pcap)/n)
SE = sqrt(0.269 * (1 - 0.269)/632) = 0.0176

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, Zc = Z(α/2) = 1.64

Margin of Error, ME = zc * SE
ME = 1.64 * 0.0176
ME = 0.0289

CI = (pcap - z*SE, pcap + z*SE)
CI = (0.269 - 1.64 * 0.0176 , 0.269 + 1.64 * 0.0176)
CI = (0.24 , 0.298)